### v1

# The Science of Data Collection: Insights from Surveys can Improve Machine Learning Models

Abstract
Whether future AI models make the world safer or less safe for humans rests in part on our ability to efficiently collect accurate data from people about what they want the models to do. However, collecting high quality data is difficult, and most AI/ML researchers are not trained in data collection methods. The growing emphasis on data-centric AI highlights the potential of data to enhance model performance. It also reveals an opportunity to gain insights from survey methodology, the science of collecting high-quality survey data.
In this position paper, we summarize lessons from the survey methodology literature and discuss how they can improve the quality of training and feedback data, which in turn improve model performance. Based on the cognitive response process model, we formulate specific hypotheses about the aspects of label collection that may impact training data quality. We also suggest collaborative research ideas into how possible biases in data collection can be mitigated, making models more accurate and human-centric.
1 Introduction
AI and ML practitioners are not the only researchers who depend on data collected from humans to achieve their aims. Social scientists have long relied on survey data collected from human subjects to quantify the population, understand public opinion, and test hypotheses about human behavior. The methods used to collect survey data have been extensively studied and refined by researchers in the field of survey methodology, which draws on social and cognitive psychology to develop theories about how humans understand, process and respond to questions in surveys (Groves et al., 2009).
Because human-labeled data is central to all stages of the AI pipeline (Plank, 2022; Mazumder et al., 2023), it is surprising that so little research in the AI literature has used social science, and survey methodology in particular, to understand the actions and motivations of the humans behind the data generating process. Insights from social science will be critical to the development of more trustworthy and human-centric models: “if we want to train AI to do what humans want, we need to study humans” (Irving & Askell, 2019).
Collecting high quality data is difficult, as decades of research in survey methodology and recent high-profile failures in opinion polling (Sturgis et al., 2016; Kennedy et al., 2017; Clinton et al., 2021) demonstrate. We worry that many researchers collecting data to train, fine-tune, or reinforce AI and ML models are not trained in data collection. A recent paper lamented: “everyone wants to do the model work, not the data work” (Sambasivan et al., 2021). This paper introduces AI researchers to the community of scientists who want to do the data work and their insights into how to collect high quality data.
We first make the case that label collection is similar to survey data collection. Then, we explore what we learn when we think of labelers as survey respondents (Section 3) and as interviewers (Section 4). In Section 5, we join the call for greater transparency in label collection methods, offering lessons from surveys and statistical methods. Throughout the paper, we use the terms labels (and labelers) to refer generally to ML annotations, such as image bounding boxes, natural language understanding labels, model evaluation data, human feedback for reinforcement learning, and other types of training data (and data collectors).
2 How Labeling is Like a Survey
Figure 1: Example of Labeling Interface for InstructGPT (from Ouyang et al., 2022)
Surveys can be self-administered, like a web or a paper-and-pencil survey, or interviewer administered, like a telephone or face-to-face survey. A web survey presents a series of questions with, most often, closed answer choices. Superficially, the process of labeling observations for ML models often looks like a web survey. Labelers see one or more prompts and associated answer choices (Figure 1). As in surveys, the task can be factual (is there a bicycle in this photo?) or opinion based (which of these responses to the prompt is most helpful?). From this perspective, the labelers are survey respondents.
However, from a different perspective, labelers are interviewers. In a survey, respondents answer each question only once, but interviewers ask the same question dozens of times. Subtle differences in the way interviews ask questions introduce interviewer effects (Groves, 1989; Schnell & Kreuter, 2005). Like interviewers, labelers respond to a similar prompt many times, and subtle differences in how they label lead to labeler effects (Geva et al., 2019). In addition, labelers, like interviewers, are provided with training and feedback (Pustejovsky & Stubbs, 2012; Fort, 2016) and are usually compensated for their labeling work.
Of course, label collection differs from survey data collection in important ways. Surveys usually ask people about themselves: their opinions, behaviors, characteristics. Labeling tasks more often involve passing judgment on an object outside of oneself (images, product reviews). However, the human feedback used in reinforcement learning may be more like a survey question: “which of the following responses is most relevant to the prompt?” Surveys ask many different questions and end when the questions run out; labeling tasks often consist of similar repeated observations, and labelers can choose when to stop.
The goal of surveys and labeling tasks is also different. Surveys ask questions of a sample of selected persons to make inference about the population. For example, say we wished to know what proportion of the US adult population does not have health insurance. We select a sample of US adults and ask them whether they have health insurance. For each respondent, we have a yes or no response to the question. [^1] We are ignoring “I don’t know” and other types of responses. The proportion of respondents without insurance in the sample is an estimate of the proportion of the population without insurance. A survey often asks many questions on related topics. Our example survey might also ask “when was the last time you visited a health care professional?” and “have you received a flu vaccine this year?” Results are often reported at the question level although relationships between questions are also of interest: for example, are people with health insurance more likely to get a flu vaccine? The quality of the analysis rests on collecting correct responses from a representative sample of US adults.
The goal in labeling work is not to estimate the percentage of the population that finds a given post offensive or detects a vehicle in a given image. [^2] Many studies collect only one or a few labels of each instance, or collapse multiple labels per instance, and thus cannot capture the perspective and diversity of a population, despite research showing that this diversity improves models (Basile et al., 2021; Sap et al., 2022) Instead, the goal in training is to learn patterns from the labeled data to predict labels for unseen observations. Thus, while it is not as important that labels reflect the population’s views at an instance level, across instances, the views of the population should be represented Mitchell et al. (2019).
3 Labelers as Respondents
We first consider what we can learn from the survey methods literature when we consider labelers as respondents.
3.1 Relevant Social Science Theories
Social scientists have developed several theories about how respondents understand survey questions and formulate answers to them.
Optimal Survey Response Process
Responding to a survey question involves several cognitive steps (Tourangeau et al., 2000; Tourangeau, 2018). The simplest version of the cognitive response process model has four steps:
  1. 1.
    Comprehension: Understand the question and the response options
  2. 2.
    Retrieval: Search memory for relevant information
  3. 3.
    Integration: Integrate the retrieved information to form an answer to the question
  4. 4.
    Mapping: Map that answer onto the provided answer choices
Ideally, a respondent proceeds through each step in order. However, at each step, the respondent can choose to repeat an earlier step. For example, considering the response options in the Mapping step may change the interpretation of the question (Comprehension) or trigger additional relevant information to come to mind (Retrieval)).
The cognitive response process model exposes why respondents sometimes give incorrect answers. At the Comprehension step, they may fail to understand the question or some of the words it uses. They may have a different understanding of some of the words than those who wrote the question. At the second step, respondents may fail to retrieve all relevant information. Some information may have been forgotten. At the third step, respondents may fail to put in the mental effort to bring together their understanding of the question with the retrieved information. At the fourth step, respondents may not find an answer choice that reflects their answer or they may edit the true answer to avoid revealing sensitive information.
Deviations from Optimal Response Process
The full survey response process outlined above places a cognitive burden on respondents. Many resort to taking shortcuts, an approach called satisficing (Krosnick, 1991; Krosnick et al., 1996). For example, respondents may retrieve only the most recent relevant information from memory (recency bias) or choose the first reasonably correct answer choice.
Satisficing relates to the cognitive miser theory in psychology, which holds that people try to minimize cognitive effort (Fiske & Taylor, 1991; Kahneman, 2011). As predicted by satisficing theory, eye-tracking studies show that respondents do not read all options in multi-choice questions (Galesic et al., 2008), and shortcuts are more common as survey length increases (Galesic & Bosnjak, 2009). Respondents in survey tasks tend not to read provided instructions Brosnan et al. (2019) or click on provided definitions (Peytchev et al., 2010), especially when they think they know the concept that is asked about Roger et al. (2006).
Acquiescence, the tendency to say “yes” to yes/no questions, regardless of content, is another shortcut response technique (Knowles & Condon, 1999). Some respondents even deliberately give incorrect answers to later questions to reduce the burden of a survey, a phenomenon called motivated misreporting (Kreuter et al., 2011; Tourangeau et al., 2012; Eckman et al., 2014).
Context Effects
Context effects concern how human perception is influenced by information perceived previously (Tversky & Kahneman, 1974; Strack, 1992). For example, a very tall person can make others seem shorter: a contrast effect. An unethical politician can make other politicians seem less ethical: an assimilation effect (Bless & Schwarz, 2010).
In surveys, attitudes and opinions are especially vulnerable to context effects, because respondents do not always have well-formed, fixed opinions that they retrieve from memory. Instead, they form opinions when asked for them, and this process can be shaped by context clues in the sentence, the response options, the instrument, or the previous questions (Zaller & Feldman, 1992; Schwarz, 2007).
3.2 Hypotheses about Label Quality
These theories about how respondents answer questions lead us to several hypotheses about the properties of the labeling task that may impact training data quality. The ML literature has investigated some of these hypotheses, but despite the isolated efforts mentioned below, there exist fundamental research gaps.
Observation Order
Theories about context effects suggest that the order in which instances or observations are presented influences the labels assigned. If a contrast effect is present, a very hateful social media post would make later posts seem less hateful than they otherwise would (for preliminary evidence of this phenomenon, see Beck et al., 2022). An order effect could also arise if labelers change their behavior over time. As they gain experience, they might become more accurate and conscientious, as suggested by Lee et al. (2022). Alternatively, they might become bored and engage in more satisficing over time, as suggested by the survey literature (Galesic & Bosnjak, 2009). Many research questions thus remain open, for instance, how and to what extent observation order impacts training data, what type of tasks are most impacted by order effects, and the implications of order effects for active learning and similar label collection approaches.
Multiple Labels
Often ML researchers want to collect multiple labels about an observation from one labeler: for example, whether an image contains a cat, a dog, a person, or a vehicle. We can ask labelers to provide all labels at once, as in the left panel of Figure LABEL:fig:multi, or we can ask one or more labelers to provide each label separately, as in the right panel (and also in Figure 1).
The choice between these two approaches echos the choice in surveys between select-all-that-apply questions and a series of yes/no questions (Figure LABEL:fig:selectall). As predicted by the survey response model, the yes/no format collects better data because it encourages respondents to process each option separately (Smyth, 2006; Pew Research Center, 2019). The select-all-that-apply approach is vulnerable to satisficing: respondents pick the first one or two reasonable options and fail to think deeply or even look at later options (Galesic et al., 2008). However, the yes/no format can encourage acquiescence (Smyth, 2006).
This finding also holds for labeling: when collecting two labels of the same observation, splitting the collection across two screens (the left panel of Figure LABEL:fig:multi) results in better performing models (Kern et al., 2023). This result is a clear example of how findings in the survey literature translate to the labeling task and improve the quality of training data.
Don’t Know Option
The inclusion of ‘don’t know’ or ‘no opinion’ responses in surveys has been debated for years. Some researchers believe these options offer respondents an easy way to satisfice: rather than thinking about the issue and forming an opinion, respondents can simply choose the ‘don’t know’ option. Others believe that having no opinion on a given topic is a valid response and that forcing respondents to provide an opinion when they don’t have one reduces data quality (Schuman & Presser, 1996).
Many labeling tasks do not include a ‘skip’ or ‘don’t know’ option. Labelers must provide a label even when they are not certain. When a recent experiment provided a ‘don’t know’ option to half of the labelers, fewer than three percent chose it, and the overall distribution of the labels was not impacted (Beck et al., 2022). Another recent study in NLP collected ‘uncertain’ flags from labelers for a relation extraction task across different text genres Bassignana & Plank (2022). Some genres were more challenging to label, and the classifier also struggled with these items. These preliminary results suggest that giving labelers the option to indicate uncertainty or lack of knowledge provides helpful additional information and does not encourage satisficing in labeling.
Pre-labeling
Pre-labeling involves displaying a suggested label, bounding box, or similar and asking the labeler if it is correct. If the labeler indicates the label is not correct, they are asked to provide the correct label. Pre-labeling is more efficient than labeling without suggested labels (Lingren et al., 2013; Skeppstedt et al., 2016; South et al., 2014). However, labelers may become too trusting of model suggestions and fail to correct errors (Dietvorst et al., 2015; Logg, 2017; Berzak et al., 2016), a phenomenon called anchoring bias or automation bias (Mosier & Skitka, 1999).
In the survey field, we find that providing a pre-filled response that respondents or interviewers should update leads to underreporting of errors of both omission and commission. For example, when respondents are reminded of their answer in a previous survey wave, they tend to report that the answer still applies rather than providing an updated response (Jäckle & Eckman, 2019).
Previous literature on labeling has explored anchoring bias (Lingren et al., 2013; Skeppstedt et al., 2016; South et al., 2014) but has not leveraged social science to find the factors that make make the effect weaker or stronger. The social science literature suggests several hypotheses about the mechanisms behind anchoring bias, such as incentives (Cialdini, 2009), belief in authority (Asch, 2016; Cialdini, 2009), or reliance on heuristics (Cialdini, 2009; Norman, 2007; Kahneman, 2011). Testing these theories experimentally would help design tasks that capture the efficiency of pre-labeling with lower risk of anchoring bias.
Overreliance on Examples
Biases similar to anchoring from pre-labeling can evolve through examples. Survey questions often give examples of the things respondents should consider when they formulate their responses. Examples improve response accuracy when they remind respondents to include items they might otherwise leave out, because they have forgotten or were unsure whether to include them. However, when the examples include only frequent items, respondents tend to leave out infrequent items more often than they did before the examples were provided (Tourangeau et al., 2014).
Labeling instructions or guidelines also often include examples. As in surveys, labelers at times rely too heavily on these examples as they label. This instruction bias can lead to overestimation of model performance (Parmar et al., 2023).
3.3 Mitigation Measures and Future Research
The survey methods literature suggests several approaches to minimize the effects introduced in Section 3.2.
Instrument testing
Many surveys spend weeks or even months drafting, testing, and revising questions and response options to arrive at language that is understood similarly by most members of the population, a process called cognitive interviewing (Willis, 2004). They then launch the survey with a small group of respondents to assess response time, don’t know rates, and response rates. Such testing could improve the instructions and prompts given to labelers.
Retain paradata
Many surveys capture process data, called paradata, during the survey, such as the time spent on each screen, the device used, even mouse movements (Kreuter, 2013; Horwitz et al., 2017). Paradata can help identify satisficing respondents and low quality data. While paradata allows for additional insights, collecting such data may raise additional privacy and ethics concerns. For a discussion of the privacy and ethics of paradata collection in surveys, see Couper & Singer (2012); Kunz et al. (2020); Henninger et al. (2023).
Feedback to Labelers
Label collection instruments could experiment with prompts to encourage labelers not to engage in satisficing. Respondents who pick many ‘don’t know’ answers or repeatedly choose the same response option can receive reminders about the importance of the task. Those who click through screens more quickly than they could possibly read questions can receive prompts to slow down and read carefully. In surveys, feedback on speeding successfully slowed respondents and did not lead to early terminations of the survey (Conrad et al., 2017).
Limit observations per labeler
Many labeling tasks allow labelers to continue as long as they like or until there are no more observations to label. However, satisficing may be more likely as the task continues, as found in surveys (Galesic et al., 2008). Thus, ML and AI researchers may wish to limit labelers to a maximum number of observations. However, such a limit may not make sense for tasks with high training costs.
Test observations
In surveys, instructed response items, such as “Choose yellow below” can help identify respondents who speed or provide low quality responses (Gummer et al., 2018). Embedding observations with known labels in a labeling task can similarly help identify satisficing labelers (Nie et al., 2020). In NLP tasks, it is common to qualify only workers who pass an initial quiz or perform well on inserted test observations Nangia et al. (2021).
We recommend future work to more comprehensively test these and related approaches.
4 Labelers as Interviewers
Thinking of labelers as interviewers rather than respondents leads to additional insights.
4.1 Interviewer effects in surveys
When survey data are collected by interviewers, as in face-to-face or phone surveys, the characteristics and behavior of the interviewer, and the interaction of these with the respondents and the questions, impact the data. For example, interviewer age, gender, and race can influence respondents’ answers (see West & Blom, 2016, for a detailed review).
The statistical model for these effects assumes that each interviewer $k$ biases the responses they collect slightly. The response to question $Y$ collected from respondent $i$ is a function of the true response and the interviewer level error term, $\delta_{k}$.
$Y^{measured}_{i,k}$ $=f(Y^{true}_{i},\delta_{k})$
$\delta_{k}$ $\sim N(0,\Delta)$
When a survey uses only a few interviewers, the individual effects can introduce bias into estimates. When there are many interviewers, the individual effects cancel out. For this reason, it is best practice in surveys to use many interviewers and make sure no interviewer dominates the data set (Kish, 1962; Groves, 1989).
Questions eliciting respondents’ opinions have larger interview effects than those about factual matters (that is, a larger $\Delta$). The explanation is the same as that cited above to explain why opinion questions are more susceptible to context effects: respondents often form opinions when asked for them, rather than drawing on pre-formed opinions. When they do so, they are impacted by the interviewer’s appearance, tone, demographics, personality, etc. (Schnell & Kreuter, 2005; West & Blom, 2016). However, interviewer effects also arise with factual and even demographic questions (Hansen et al., 1960).
4.2 Labeler Effects
In a similar way, each labeler $k$ may have a small effect on the labels they assign, $L^{collected}$, which we might call $\lambda_{k}$.
$L^{collected}_{i,k}$ $=f(L^{true}_{i},\lambda_{k})$
$\lambda_{k}$ $\sim N(0,\Lambda)$
Although the ML literature has not conceptualized labeler effects in this way, we find evidence in the literature that these $\lambda$ values are non-zero and correlate with labeler characteristics. For example, labeler age and education level influence whether they perceive comments on Wikipedia entries as attacks (Al Kuwatly et al., 2020). Conservative labelers are less likely to flag anti-Black language (Sap et al., 2022). [^3] To our knowledge, most evidence of labeler effects comes from the NLP literature. However, our experience with interviewer effects across all types of questions leads us to believe that labeler effects also occur in image labeling, human feedback for reinforcement learning, and all types of label collection. As with surveys, the labeler effects are likely smaller for some tasks than others. More objective tasks that most labeler agree on may have small effects. Tasks like hate speech annotation probably have larger $\lambda$ values. These labeler effects also impact models: training a model on subsets of labelers with shared characteristics leads to different predictions (Al Kuwatly et al., 2020; Perikleous et al., 2022).
Just as with interviewers, it is preferable to have many labelers each with a reasonable number of labels rather than a few labelers with many labels. The argument is the same: When many labelers contribute to a data set, their individual effects cancel out and do not result in systematic bias. As discussed above, studies may find it impractical to have many labelers.
Although many consider the $\lambda$s noise, as we do in surveys, others have argued that they are in fact information and can improve model performance. Embracing label variation may make models more trustworthy (Aroyo & Welty, 2015; Plank, 2022).
4.3 Who Works as Labelers
Training models on labels that reflect the views of the population will increase the inclusiveness and trustworthiness of models. However, our experience with survey data collection leads us to worry about how well labelers represent the population.
Large-scale annotation tasks, such as the reCAPTCHA tests, may collect labels from a representative slice of the population. However, the crowdworkers used in many label collection tasks are members of large panels such as Appen, Upwork, Scale AI, Prolific, or MTurk and do not reflect the US or world population. The workers who labeled data to fine-tune InstructGPT were members of Scale AI and Upwork; 22% were Filipino and 22% were Bangladeshi (Ouyang et al., 2022, Appendix B3). MTurk members are younger, lower income and less likely to live in the South than the US population (Berinsky et al., 2012).
In surveys, selection bias [^4] In the survey literature, the preferred terms is nonresponse bias. We use the term selection bias here because it is more general and more suited to the labeling task. arises when the propensity to take part in a survey is correlated with the characteristics measured in a survey. Return to the example in Section 2, a survey to estimate the proportion of US adults without health insurance. If we distributed the survey invitations in doctors’ waiting rooms, the sample proportion would overestimate the population proportion. Those in waiting rooms are more likely to receive the survey invitation and are also more likely to have health insurance. The propensity to take part is correlated with what the survey measures, leading to selection bias in the estimate of health insurance coverage.
Just as in surveys, selection bias arises in training data labels if the propensity to engage in the labeling task is correlated with the propensity to assign a given label. We expect that this correlation is non-zero for many tasks, for two reasons. The first reason we expect a correlation is that labeler characteristics likely influence both propensities, as shown in Figure 2. On the right side of the figure, labelers’ characteristics influence their decision to engage in labeling work, as discussed earlier in this subsection. On the left side of the figure, labelers’ characteristics influence the labels they assign, as discussed in the previous subsection on Labeler Effects (Section 4.2). If the characteristics that influence one side of the figure are the same as, or correlate with, the characteristics that influence on the other side of the figure, the two propensities at the bottom will be correlated, leading to selection bias.
The second reason we expect a correlation is that labeling tasks are largely opt-in and labelers can see the type of work they will be involved in before deciding to participate. For example, someone with strong views on bicycle safety may be more likely to see bike lane violations where others do not and, for that reason, be more likely to choose to label a data set of images of potential bike lane violations.
Figure 2: Labeler characteristics impact propensity to assign a given label and propensity to participate in labeling
4.4 Mitigation Measures and Future Research
To combat selection bias in labeling tasks, we need to break the correlation between the propensity to assign a given label and the propensity to participate in labeling. One way to do so is to remove the correlation between labeler characteristics and the propensity to participate in labeling by diversifying the labeler pool, collecting data from labelers with different motivations and characteristics. This approach is central to surveys: we solicit responses from a random sample of the population through appeals to public service, tokens of appreciation, and multiple reminders (see, for example, Groves & Couper, 1998). If we can collect responses from labelers who represent the population, selection bias is not a problem. Unfortunately, it is not clear that this approach will work with label collection: many people are not interested in labeling data for AI models. And some tasks require expert knowledge, such as reading radiographs or annotating low-resource languages; for these tasks, representative samples are likely not possible. (Surveys also find it increasingly difficult to recruit representative samples of respondents (de Leeuw E., 2018; Williams & Brick, 2018)).
Another way to break the correlation in Figure 2 is through statistical adjustment. The survey literature contains many statistical methods to match the characteristics of the respondents to the population and thus reduce selection bias (Bethlehem et al., 2011, for example). For example, surveys in many countries collect more responses from women than men; we use weights to equalize the contribution of women’s and men’s responses on the final estimate. Future work could test the usefulness of these weighting approaches for improving machine learning models. [^5] The tendency for some crowdworkers to give false answers to demographic questions to protect their privacy, reported by Huang et al. (2023), will complicate any weighting approach. However, these statistical adjustments can work only if we know and measure the labeler characteristics that drive the relationships in Figure 2. Thus we need a better understanding of what motivates people to work as labelers and what types of tasks are vulnerable to labeler effects (the two paths in Figure 2), which points to another role for social science to play in improving label quality.
5 Transparency in Label Collection
The discussion above suggests that labels are sensitive to how studies design the labeling task and recruit, train, and compensate labelers, in ways not recognized in the AI/ML literature. For this reason, we call for more transparency and documentation in how labels are collected when new data sets or models are released.
The survey industry in the U.S. has embraced transparency in recent years. The American Association for Public Opinion Research launched the Transparency Initiative in 2014. Survey data collection firms that are accepted as members agree to disclose details about how survey data were collected when requested, such as sponsor, question wording, and weighting adjustments. [^6] See https://aapor.org/wp-content/uploads/2022/11/TI-Attachment-C.pdf for details. Polling companies that are members of the Transparency Initiative outperform those who are not, suggesting that willingness to disclose methods is a proxy for the quality of a firm’s data collection methods (Silver, 2023). The U.S. federal statistical agencies recently commissioned an expert report on how they can improve transparency as well (National Academies of Sciences, Engineering, and Medicine, 2022), another sign that transparency in data collection has positive externalities.
Several researchers have similarly called for transparency when releasing benchmark data sets for AI and ML models (Bender & Friedman, 2018; Hutchinson et al., 2020; Gebru et al., 2021; Chmielinski et al., 2022). However, none of these standards request the kind of data collection process information required by the Transparency Initiative. The absence of more detailed documentation of data collectinon methods may be behind the recent difficulties in replicating benchmark data sets (Recht et al., 2019). We commend Nie et al. (2020); Ouyang et al. (2022) as particularly good examples of transparency in label collection, with the latter including screenshots of the label collection instrument (Figure 1). Prabhakaran et al. (2021); Geiger et al. (2020); Baan et al. (2024) have also called for better documentation of the label collection process.
6 Outlook
Collecting data from respondents or labelers is more difficult than the AI/ML literature has recognized. Opinion questions are particularly difficult, because opinions can be shaped by wording, order, and other context effects and by the characteristics of the interviewer or the labeler. We believe that the type of human feedback needed to ensure that future models are trustworthy will likely resemble opinion collection in surveys, as in Figure 1. For this reason, greater cooperation and knowledge sharing between the AI/ ML and survey methods fields will be crucial to ensuring that the next generation of even more powerful models is human-centric, trustworthy, and fair.
Although some hope to replace human labelers with models-as-labelers (Pangakis et al., 2023), the data collection challenges described in this paper will not go away. First, because models are trained on data collected from humans, they can display many of the response biases described in this paper (Dominguez-Olmedo et al., 2023; Tjuatja et al., 2023). Second, models trained on data labeled by models exhibit unusual behavior, called Model Autophagy Disorder (Alemohammad et al., 2023). Third, for the foreseeable future, we will still want humans to create the labels that teach models how to be fair and safe. Thus, the most difficult and most important labeling must still be done by humans.
As is often the case when working across disciplines, as this paper does, language and cultural differences get in the way of information sharing. The survey methodology literature tends to focus on uncovering bias and variance and attributing it to a cause (interviewers, question wording, survey mode). AI researchers, on the other hand, often understandably focus on solutions in the form of python packages or github repos. These cultural differences between the fields can leave researchers unable to present at each other’s conferences or even have productive conversations. Nevertheless, we hope that this position paper has made the case that theory-driven social science can help AI researchers develop insights and tools that improve the quality of their data and their models. And solution-oriented thinking could also make survey data better. Perhaps future joint workshops could help the two fields share insights.
We applaud the growing interest in data-centric AI, which focuses on improving models by improving the data on which they rely rather than using the latest algorithm. We recommend that researchers developing AI models be as careful with how their data are collected as they are with their models. We also suggest reaching out to applied statistics departments and social science research groups which often focus on the science of data and may be willing to do the data work.
Acknowledgements